40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17024     Entry time: Wed Jul 20 18:07:52 2022     In reply to: 17020     Reply to this: 17037
Author: Paco 
Type: Update 
Category: BHD 
Subject: BHD MICH test 

[Paco, Yuta, JC]

We did some easy tests on the BHD readout in preparation for BHD MICH. With the arm cavities and LO beam misaligned, but the MICH aligned, we measured the transfer function from C1:LSC-DCPD_A_OUT to C1:LSC-DCPD_B_OUT to get a rough estimate of the gain balance: 1.8 * DCPD_A = DCPD_B. We then locked MICH using REFL55_Q and looked at

  • A=C1:LSC-DCPD_A_OUT
  • B=C1:LSC-DCPD_B_OUT
  • 1.8 * A - B (which we encoded using C1:LSC-PRCL_A_IN1)
  • 1.8 * A + B (which we encoded using C1:LSC-PRCL_B_IN1)

namely the DCPD BHD signals. After turning the MICH_OSC on (2000 gain @ 311.1 Hz), we took some power spectra under the following three configurations:

  1. LO misaligned, no MICH offset.
  2. LO overlap, no MICH offset.
  3. LO overlap and MICH offset.

For 1. the expectation was that since LO is misaligned and the AS port is dark, we would get no signal. In 2., however both A and B would might see some incoherent signal, but still no MICH. Finally in 3. all signals should be able to see MICH, including A-B. Attachment #1 shows the measurements 1, 2, and 3 (offset = -5.0). Then, with increasing offset values, the BHD MICH signals increased as well; discussion to follow.

Attachment 1: BHD_MICH_OSC.pdf  19 kB  Uploaded Wed Jul 20 19:17:40 2022  | Hide | Hide all
BHD_MICH_OSC.pdf
ELOG V3.1.3-